OpenStudy (anonymous):

solve for x and y: {y=5x+6;y=9x+7}

OpenStudy (anonymous):

Simplify first,; 5x + 6 - 6 = y - 6 9x + 7 - 7 = y - 7 5x = y - 6 9x = y - 7

OpenStudy (anonymous):

my teacher said to set them equal to each other o.o'

OpenStudy (anonymous):

both are equal to y means the expressions in the right hand side of both the equation are equal